NON-EXISTENCE OF A BIPLANE (121,16,2) WITH THE AUTOMORPHISM GROUP A_{5} by Slavka Pfaff, Zagreb, Croatia

Abstract. No biplane (121,16,2) has ever been found. It is shown that A_{5} cannot operate on such a design. Mathematics subject classification: 05B05 Keywords: symmetric design, biplane, automorphism group, orbital structure.